OpenStudy (anonymous):

Prove or Disprove Let a, b be integers. If a/b^2, then a/b

OpenStudy (anonymous):

Not sure what you're asking. Are you asking "if a divides b^2, then a divides b"? Or are you suggesting the fraction \[ \frac{a}{b^2} \text{or } (\frac{a}{b})^2\] In which case "If a/b^2" isn't really a statement

OpenStudy (anonymous):

lol thats what it is on a divides b^2 it thats true then a divides b. I have to prove this statement

OpenStudy (anonymous):

Ok that's what I thought. I'm sorry, I messed up the order initially. Can you think of a counterexample? What if a = 9 and b=6? Does 9|36? Yes. Does 9|6? No. Sorry, I was confusing myself while writing this, let me know if you have any additional questions.

OpenStudy (anonymous):

Thanks, yea I got another questions Let a,b,c be integers. If c > 0, then gcd(ac,bc) = cgcd(a,b). Suggestion: Let d= gcd(a,b). Show cd = gcd(ac,bc)

OpenStudy (anonymous):

If you think about this one intuitively, it should make sense. To give you a rough idea: \[ d = \text{gcd}(a, b) \] \[ \Rightarrow \exists m,n \in \mathbb{Z} \text{ such that } dm=a, \ dn = b \] Multiply both sides of each equality by c \[ c(dm)=ac, \ c(dn) = bc \] \[ m(cd) = ac, \ n(cd) = bc \] \[ \Rightarrow cd = \text{gcd} (ac,bc)\]

OpenStudy (anonymous):

Thanks